OpenStudy (anonymous):

What is the mean of this discrete random variable? That is, what is E(X), the expected value of X? X 24 26 27 32 35 39 P(x) 0.16 0.09 0.18 0.12 0.24 0.21

OpenStudy (kropot72):

\[E(X)=24\times 0.16+26\times 0.09+27\times 0.18+32\times 0.12+35\times 0.24+39\times 0.21\]